OpenStudy (dls):

A wire in the form of a circular loop of radius 10 cm lies in a plane normal to a magnetic field of 100 T.If this wire is pulled to take a square shape in the same plane in 0.1 s,find the average induced EMF in the loop.

OpenStudy (dls):

Attempt: \[\LARGE \phi = BA\] \[\LARGE \frac{d \phi}{dt}=B \frac{dA}{dt}\] For change in area.. 2 PI r= L (Lets compute in m) 2 pi /10 = L \[\LARGE L^2=\frac{4 \pi^2}{100}\] (Final area i.e square's) Initial Area=> \[\LARGE \frac{\pi}{100}\] \[\LARGE e_{induced}=100 \times \frac{\pi (4\pi-1)}{100 \times 0.1}\] \[\LARGE e_{induced}=10 \pi (4\pi-1) =>363V\] somehow wrong

ganeshie8 (ganeshie8):

how did u get this : 2 PI r= L (Lets compute in m) 2 pi /10 = L

ganeshie8 (ganeshie8):

2 PI r =perimeter of wire right ?

OpenStudy (dls):

Perimeter of circle=Perimeter of square :O ?

ganeshie8 (ganeshie8):

Is L perimeter of square ?

OpenStudy (dls):

oh L/4 ?

ganeshie8 (ganeshie8):

yeah, better call L , the side of square

ganeshie8 (ganeshie8):

then area of square = L^2

OpenStudy (dls):

L is perimeter..then L/4 is side..i.e length of square side = 2pi/40 Area of square=> \[\LARGE Area_{sq}=\frac{4\pi^2}{1600}=>\frac{\pi^2}{400}\]

ganeshie8 (ganeshie8):

okie..

ganeshie8 (ganeshie8):

hey wait

OpenStudy (dls):

Then dA/dt=> \[\LARGE \frac{dA}{dt}=\frac{\pi}{100}-\frac{\pi^2}{400}=>\frac{\pi}{100}(1-\frac{\pi}{4})\]

OpenStudy (dls):

:O

ganeshie8 (ganeshie8):

length of square side = L / 4 = 2 Pi r / 4 = 20 Pi / 4 = 5 pi

ganeshie8 (ganeshie8):

right ?

OpenStudy (dls):

that is in CM right?

OpenStudy (dls):

okay lets do in cm then, \[\LARGE \frac{dA}{dt}=100\pi-5\pi=95\pi\]

OpenStudy (dls):

\[\LARGE e_{induced}=100 \times 95\pi \times 10 =>95000\pi\] omg

ganeshie8 (ganeshie8):

\(\LARGE \frac{dA}{dt}=\frac{100\pi-5\pi}{0.1} \)

OpenStudy (dls):

into 100 wahi to likha

OpenStudy (dls):

dA/dt=950 pi

ganeshie8 (ganeshie8):

scratch that \(\LARGE \frac{dA}{dt}=\frac{100\pi-(5\pi)^2 }{0.1} \)

OpenStudy (dls):

emf=950 x 100 = 95000 pi

OpenStudy (dls):

lol nvm

OpenStudy (dls):

\[\LARGE \frac{dA}{dt}_{final}=>250\pi(4-\pi)\]

OpenStudy (dls):

\[\LARGE emf_{Final}=>100 \times 250\pi(4-\pi)=>25000 \pi(4-\pi)\]

OpenStudy (dls):

\[\Huge 67510 \times 10^{-4}=>6.75 V\] Winner!

ganeshie8 (ganeshie8):

:) u just missed /4 for L in ur first attempt

OpenStudy (dls):

sare physics scientist ab online aa rahe hai -.-

OpenStudy (dls):

sigh* yeah :|